OpenStudy (ineedhelpnow):

solve the linear system 3x-4y=8 4x+5y=10

OpenStudy (anonymous):

multiply eq 1 by 5 and eq 2 by 4 then add both eq. we get 31x=80 or, x=80/31 substitute it in either equation and solve for y.

OpenStudy (ineedhelpnow):

nvm (3x-4y=8)5 (4x=5y=-10)4 ------------- 15x-20y=40 16x+20=-40 ------------- 31x=0 -- -- 31 31 _______________ x=0 3(0)-4y=8 _______________ -4y=8 -- -- -4 -4 ------------- y=-2

OpenStudy (anonymous):

In question's second eq there is +10 not -10.

OpenStudy (ineedhelpnow):

oops! wrote it wrong the first few times. 4x+5y=-10.